OpenStudy (anonymous):

How would I graph x + y = 11 if my graph only goes 10 over and -10 down?

OpenStudy (anonymous):

Extend your graph..?

OpenStudy (anonymous):

Wouldn't you do y = -x + 11 Slope equals -1 and just graph a diagonal line down from the top?

OpenStudy (anonymous):

I can't, unfortunately.

OpenStudy (anonymous):

Hit girl, that sounds right but can you clarify a bit more? Do you know what the ordered pair would be fore this equation?

OpenStudy (anonymous):

Wouldn't you do y = -x + 11 Slope equals -1 and just graph a diagonal line down from the top? U answered your own question...:-)

OpenStudy (anonymous):

thanks estudier. I have that link and tried it. They show a graph that goes well beyond 10 and -10. It has the line graphed at 11 up and 11 over which I can't do on the graph provided.

OpenStudy (anonymous):

You haven't tried my link, then because I restricted it to -10, 10 on both axes.

OpenStudy (anonymous):

oh sorry...

OpenStudy (anonymous):

You're saying it goes 10 to the right of the y-axis, and 10 below the x-axis?

OpenStudy (anonymous):

Haha yeah... I did. I'm new to this anyway. Hope it helped somewhat =P

OpenStudy (anonymous):

Yes oncoming...I tried to slope it as 10 up and 10 over and it said that was not the right answer. The graph only goes to 10.

OpenStudy (anonymous):

Have u looked at the link now?

OpenStudy (anonymous):

Are you looking for any ordered pair? You can just make an x/y table and set some values for x to see what y is

OpenStudy (anonymous):

Yes, and it looks as though it is sloped at 10 up and 10 over. I tried that and it said it wasn't the right answer. Maybe I am not reading the graph right on the link you provided.

OpenStudy (anonymous):

If this x + y = 11 is the equation then the link I gave u shows the correct graph.

OpenStudy (anonymous):

Okay, I think I just figured out what I did wrong. thanks.